Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for May 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Friday, May 20, 1998

FOR DATES: From 7:00 a.m. Wednesday, May 13, 1998 to 7:00 a.m. Wednesday, May 20, 1998.

The Thermal Destruction Facility (TDF) was on-line for approximately 101.4 hours and processed approximately 4281 tons of contaminated soil during the period of this report. The TDF was restarted early Wednesday, May 13, 1998 after the feed screws had been replaced and the kiln and SCC were re-heated to operating temperatures. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

To date, the contractor has successfully treated approximately 43,968 tons of contaminated soil through the TDF and all ash analyzed during the operations phase has met project backfill criteria.

On two occasions during the time frame of this report, feed was stopped to perform maintenance on the acid gas scrubber system. On both occasions, Friday morning and Friday night, the contractor's TDF control room operators observed unusually high temperatures within the scrubber system. In order to perform equipment investigations within the scrubber system, feed to the kiln was stopped and a planned Thermal Relief Valve (TRV) opening was scheduled in order to lower temperatures within the stack system. After the appropriate authorities were notified, the TRV was opened and an investigation by the contractor began.

Investigations indicated that the venturri spray nozzles within the acid gas scrubber were clogged with rubber debris. Further investigations early Saturday morning revealed that the source of the rubber material was from damaged or deteriorated acid gas scrubber pump seals. The pump was replaced, the scrubber system was flushed and cleaned by the contractor, and the system was back on-line by 0557 hours, Sunday, May 17, 1998.

Construction activities during this period included wastewater management, excavation of soils within the exclusion zone, rock washing and decontamination of inorganic debris for backfilling, excavation and handling of drums, placement of Concover on the stockpile and odor controlling foam on the excavation face within the exclusion zone, and continuation of off-site and perimeter air monitoring.

During the time frame of this report, 47 drums were excavated, sampled, tagged, and over packed. The over packed drums are currently in temporary storage within the exclusion zone. The contractor is awaiting finger print lab analyses (characterization) of the drums in order to safely group or bulk the drums prior to shipment off-site for final disposition.

Related site activities during this period included:

* Weekly progress meeting held on Tuesday, May 19, 1998. * Dioxin stack testing performed and completed during the time frame of this report.

Major activities for the next week include: Operating the TDF, soil feed preparation, off-site and perimeter air monitoring, WWTP operations, ash handling, rock & inorganic debris washing, and drum handling, labeling, fingerprinting, and shipment off-site.

PREPARED BY: Steve Gibson & David Modricker, USACE

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.